A pension fund must pay $10,000 at the end of each of the next five years. Find the Present Value and Duration for this set of payments using a discounted rate of 0.06 or 6 %.
1
Expert's answer
2018-08-29T10:06:09-0400
If a pension fund must pay $10,000 at the end of each of the next five years, then the Present Value for this set of payments using a discounted rate of 0.06 or 6 % is:
PV = 10,000/1.06 + 10,000/1.06^2 + 10,000/1.06^3 + 10,000/1.06^4 + 10,000/1.06^5 = 42,123.64.
Duration = 42,123.64/10,000 = 4.21.